You paid $965 for a corporate bond that has a 7.36% coupon rate. What is the current yield?
Answer #1

Given

Coupon rate = 7.36%

Coupon amount = Coupon rate * face value of bond (assuming bond FV as 1000$) = 1000*7.36/100 = $ 73.6

Current Yield = Coupon amount/Current market price of bond

= (73.6/965)*100

Current yield = 7.6269% or 7.627%